OpenStudy (anonymous):

The number of computers sold by BCC depends on the dollar amount, x, that they spend on advertising. How many computers will they sell by spending $100,000 on advertising? Round to the nearest whole number and do not include units in your answer. N(x) = 100 + 20*ln(0.25x)

OpenStudy (anonymous):

replace x by 100,000 and see what you get: \[N(100000)=100+20\times\ln(.25\times 100000)\]

OpenStudy (anonymous):

\[100+20\times \ln(25000)\]

OpenStudy (anonymous):

now use a calculator. i get 302.5 rounded